A Special rubber ball is dropped from a height of 128 meters . Each time the ball bounces to half its previous height. After 1 bounce the ball reaches a height of 64 meters.
a. How high did the ball reach after the 7th bounce.
b.How high does the ball get after the 9th bounce

Answer by windsolace(8)

The series identified is a geometric progression with common ratio of half.
Formula for geometric progression:
Nth term of series = ar^(n-1), where a = first term value, r = common ratio
a) 7th term of the series = 64(1/2)^(7-1) = 64(1/2)^6
= 64(1/64)
= 1
The ball reached a height of 1m after the 7th bounce

b) 9th term of series = 64(1/2)^8
= 64/256
= 0.250
The ball reached a height of 0.25m (25cm) after the 9th bounce.